Assume that the readings at freezing on a batch of thermometers are normally distributed with a mean of 0°C and a standard deviation of 1.00°C. A single thermometer is randomly selected and tested. Find P55, the 55-percentile. This is the temperature reading separating the bottom 55% from the top 45%.
Answer #1

solution

Using standard normal table,

P(Z < z) = 55%

= P(Z < z) = 0.55  

= P(Z <0.13 ) = 0.55

z =0.13 ( Using standard normal z table,)

Using z-score formula  

x= z * +

x=0.13 *1.00+0

x= 0.13